A wheel with a weight of 390 N comes off a moving truck and rolls without slipping along a highway. At the bottom of a hill it is rotating at an angular velocity of 24.0 rad/s . The radius of the wheel is 0.555 m and its moment of inertia about its rotation axis is 0.800 MR2. Friction does work on the wheel as it rolls up the hill to a stop, at a height of h above the bottom of the hill; this work has a magnitude of 3490 J .

1.) Calculate h. Use 9.81 m/s2 for the acceleration due to gravity.
Answer #1

Mass of wheel = 390/g = 39.76 kg.

Velocity at bottom = ωr = 24*.55= 13.32m/s

Total KE at the bottom of the hill = translational KE of CM plus rotational KE about the

CM. = (1/2)mv^2 + (1/2) I ω^2 = (1/2)mv^2 + (1/2)(.8)mv^2 = 0.9mv^2 = 6348.9 Joules.

Subtracting the magnitude of the work done by 'friction' you are left with 6348.9 - 3490 =2858.9joules that will be converted into gravitational potential energy:

2858.9= mgh = 39.76*9.8 h or h = 7.34meters in height.